Please help with the following problem. Thanks so much for your help! I really appreicate it.
It is approximately 480 miles from New York to Virginia. On a good day, a driver can average approximately 60 miles per hour. Answer the following questions using the information above.
a) How far have you traveled after 3 hours?
b) How far have you traveled after 4 hours?
c) How far have you traveled after t hours? Explain by writing a linear function that expresses the distance traveled, d, as a function of time, t.
d) How far will you have left to travel to reach Virginia after you have traveled 3 hours?
e) How far will you have left to travel to reach Virginia after you have traveled 4 hours?
f) How far will you have left to travel to reach Virginia after you have traveled t hours? Explain by writing a linear function that expresses the distance to be traveled to reach Virginia, s, as a function of time, t.

Answer by josmiceli(19441)

(a) 60 mi/hr)(3 hr) = 180 mi
(b) 60 mi/hr)(4 hr) = 240 mi
(c) d = 60t
(d) 480 - 60*3 = 300 mi
(e) 480 - 60*4 = 240 mi
(f) s = 480 - 60t